OpenStudy (anonymous):

Find and solve a system of linear equations to answer the question below. Joanne buys one kilogram of French Roast and one kilogram of Boxer Blend from the Java Hut. Jacob buys two kilograms of French Roast and three kilograms of Boxer Blend. If Joanne spends $14 for her purchase, and Jacob spends $36 for his, what is the price of one kilogram of Boxer Blend coffee?

OpenStudy (radar):

Let x = the cost of Boxer Blend coffee per kg Then $14-x cost of French Roast per kg 2(14-x) +3(x)=36 28-2x+3x=36 x=8 or 8$ per kg of Boxer Blend 14-8=6 or $6 per kg of French roast